Let X denote the amount of space occupied by an article placed in a 1-ft3 packing container. The pdf of X is below.

f(x) =

72x7(1 − x)      0 < x < 1
0 otherwise

(a) Graph the pdf.



Obtain the cdf of X.

F(x) =
     0 x < 0
0 ≤ x ≤ 1
     1 x > 1



(a) Using the cdf from (a), what is P(0.3 < X ≤ 0.6)? (Round your answer to four decimal places.)


(b)What is P(0.3 ≤ X ≤ 0.6)? (Round your answer to four decimal places.)


(c) What is the 75th percentile of the distribution? (Round your answer to four decimal places.)


() Compute E(X) and σX. (Round your answers to four decimal places.)

E(X) =
σX =


(f) What is the probability that X is more than 1 standard deviation from its mean value? (Round your answer to four decimal places.)
Answer #1
this is beta distribution with parameter α=8 and β=2

a)

F(x)= 9x^8-8x^9


b)

P(0.3<X<=0.6)= 0.0701
P(0.3<=X<=0.6)= 0.0701

c)

75 th percentile: = 0.8928

d)

E(X)=α/(α+β)= 0.8000
σx=√(αβ/((α+β)2*(α+β+1))= 0.1206

f)

P(X<E(X)+σx)+P(X>E(X)+σx)= 0.3182